# references.bib

### v1

@@ 4 Nonlinear subwavelength resonances
It is important to note that the techniques in this section are not limited to $g(u)=\lvert\frac{\partial u}{\partial t}\rvert^{2}\frac{\partial u}{\partial t}$, but can be used for any $g(u)$ that satisfies that its Fourier transform is of the order $O(\omega^{2})$.
@@
$\left\{\begin{array}[]{rll}\displaystyle\Delta u+\frac{\omega^{2}}{{c_{r}}^{2}}u&=i\beta\lvert\omega\rvert^{2}\omega\lvert u\rvert^{2}u&\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial u}{\partial\nu}&=\delta\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u]&\text{on }\partial\operatorname{{\mathcal{D}}},\end{array}\right.$
@@
4.1 Variational formulation
The following 2-form will play the role of $a_{\omega,\delta}$ in the nonlinear setting.
@@ 4.1 Variational formulation
Suppose that the smooth, bounded domain $\operatorname{{\mathcal{D}}}\in\operatorname{{\mathbb{R}^{3}}}3$ has well-separated connected components $B_{1},\ldots,B_{N}$ with connected boundaries. Define the $2$-form $\tilde{a}_{\omega,\delta}:H^{1}(\operatorname{{\mathcal{D}}})\times H^{1}(\operatorname{{\mathcal{D}}})\rightarrow\mathbb{C}$ associated $\operatorname{{\mathcal{D}}}$ by
@@
Recall here that $H^{1}(\operatorname{{\mathcal{D}}})\subset L^{4}(\operatorname{{\mathcal{D}}})$. With the form $\tilde{a}_{\omega,\delta}$ it is possible to characterize the resonances of the nonlinear system (6) in a similar fashion as in the linear case in Theorem 3.5 and Corollary 3.6. To this end, we will assume that there exists a neighborhood $U$ of $0\in\mathbb{R}$ and a neighborhood $V$ of $0\in\mathbb{C}$ such that for all $\delta\in U$ and for all $\omega\in V$ and for any $f\in H^{1}(\operatorname{{\mathcal{D}}})^{*}$ the following variational problem:
@@
has a unique solution $u_{f}(\omega,\delta)$. With this property at hand, the following variational characterization of small resonances of the nonlinear system (6) is possible.
@@
satisfies
$\int_{B_{i}}u_{f}\operatorname{dvol}=a_{i},\quad\forall i\in\{a_{1},\ldots,a_{N}\}.$ (16)
@@
4.2 Characterization of the resonances
In this subsection, we will determine the solution of the variational problem (15) and then derive asymptotics for subwavelength resonances of the nonlinear system (6). To this end, we consider the PDE formulation of the variational problem (15)
@@ 4.2 Characterization of the resonances
This proposition allows to derive the following fundamental theorem of nonlinear subwavelength physics by deriving the asymptotics of the nonlinear subwavelength resonances in terms of $\delta$.
@@
Let $\{q_{0},v_{2},\ldots,v_{N}\}$ be a basis of eigenvectors of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$, let $V$ be as in Definition 2.3 and denote by $\Pi:\mathbb{C}^{N}\rightarrow\mathbb{C}q_{0}$ the projection along $\operatorname{span}(v_{2},\ldots,v_{N})$ onto $\mathbb{C}q_{0}$. Then the subwavelength resonances $\omega(\delta)=\omega_{0}\sqrt{\delta}+\omega_{1}\delta+O(\delta^{\frac{3}{2}})$ of (6) associated to $q_{0}$ satisfies the following asymptotics
$\omega_{0}\text{ is an eigenvector of }\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\text{ with eigenvector }q_{0},$
$\omega_{1}=\frac{q_{0}\cdot\Pi\left[\frac{-i}{4\pi{c_{0}}}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}})q_{0}-\lvert\omega_{0}\rvert^{2}i\beta{c_{r}}^{2}(V^{vol})^{2}\lvert q_{0}\rvert^{2}q_{0}\right]}{2\lVert q_{0}\rVert^{2}},$
where $\lvert q_{0}\rvert^{2}q_{0}$ is to be understood to act component-wise.
@@
up to an error of order $(\omega(\delta)^{2}+\delta)^{2}$, where $\lvert q(\delta)\rvert^{2}q(\delta)$ is to be understood component-wise. Making the Ansatz $\omega(\delta)=\omega_{0}\sqrt{\delta}+\omega_{1}\delta+O(\delta^{\frac{3}{2}})$ and $q(\delta)=q_{0}+q_{1}\sqrt{\delta}+O(\delta)$ and inserting it into the above equation leads to the following cascade of equations:
@@
It then follows that $\omega_{0}$ is an eigenvector of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ with eigenvector $q_{0}$ and $\omega_{1}$ is given by
@@
Note that (18) has some similarities with the discrete Gross-Pitaevskii and Aubry-André-Harper models, which are used as discrete approximations of nonlinear Schrödinger equations (with cubic nonlinearities) in quantum mechanics; see, e.g., [soliton3, soliton4, soliton6]. Such quantum mechanical models are typically based on tight-binding approximations and include only nearest-neighbour interaction terms (long-range interactions are neglected). They correspond to (18) with $\operatorname{cap}^{gen}$ being tridiagonal [flach]. We refer to [weinstein1, weinstein2, weinstein3, localization, pankov] for their analysis.
@@
As in [cochlea], if we consider the scattering of an incident wave $u_{in}$ by a system of nonlinear subwavelength resonators, then based on Theorem 4.5, we can approximate the scattered field $u-u_{in}$ by a few subwavelength modes involving only the nonlinear subwavelength resonances and their associated eigenmodes, so as to retain the physically derived coupling between resonators.
As already mentioned at the beginning of this section, the above procedure is not limited to $g(u)=\lvert\frac{\partial u}{\partial t}\rvert^{2}\frac{\partial u}{\partial t}$ but solely relies on the property that the “Fourier transform” of $g(u)$ is of order $O(\omega^{3})$ as $\omega\rightarrow 0$. The following remark is an illustration of this fact if the order is $O(\omega^{2})$ as $\omega\rightarrow 0$.
@@
If one changes the nonlinearity $i\beta\lvert\omega\rvert^{2}\omega\lvert u\rvert^{2}u$ to $i\beta\omega^{2}\lvert u\rvert^{2}u$, the nonlinearity is of order $\omega^{2}$ thus influencing the leading-order behavior of the subwavelength resonances. To be precise, the cascading system of equations from the proof of Proposition 4.4 is then given by
$-\Delta u_{p,k,l}+\sum_{i=1}^{N}\left(\int_{B_{i}}u_{p,k,l}\operatorname{dvol}\right)1_{B_{i}}$ $=\frac{1}{{c_{r}}^{2}}u_{p-2,k,l}+i\beta\left(\lvert u_{f}\rvert^{2}u_{f}\right)_{p-2,k,l}+\sum_{j=1}^{N}a_{j}1_{B_{j}}\delta_{p=0}\delta_{k=0}\delta_{l=0},$
@@
$\begin{split}u_{f}(\omega,\delta)=\sum_{j=1}^{N}\frac{a_{j}}{\lvert B_{j}\rvert}1_{B_{j}}+\omega^{2}\left(\sum_{j=1}^{N}\frac{a_{j}}{{c_{r}}^{2}\lvert B_{j}\rvert^{2}}1_{B_{j}}+i\beta\sum_{j=1}^{N}\frac{\lvert a_{j}\rvert^{2}a_{j}}{\lvert B_{j}\rvert^{4}}1_{B_{j}}\right)\quad\quad\quad\quad\quad\quad\\ +\delta\sum_{j=1}^{N}a_{j}\left(-\sum_{l=1}^{N}\frac{\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}}{\lvert B_{j}\rvert\lvert B_{l}\rvert^{2}}1_{B_{l}}+\tilde{u}^{j}_{0,1}\right)\quad\quad\quad\quad\\ +\omega\delta\sum_{j=1}^{N}a_{j}\left(\sum_{l=1}^{N}\left(\frac{-i}{4\pi\lvert B_{j}\rvert\lvert B_{l}\rvert^{2}{c_{0}}}(\operatorname{cap}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}}))_{l,j}\right)1_{B_{l}}+\tilde{u}^{j}_{1,1}\right)+O((\omega^{2}+\delta)^{2}).\end{split}$
In that case, the subwavelength resonances $\omega(\delta)=\omega_{0}\sqrt{\delta}+O(\delta)$ are characterized by the nonlinear eigenvalue problem
$\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})q_{0}-\omega_{0}^{2}\left(q_{0}+i\beta{c_{r}}^{2}(V^{vol})^{2}\lvert q_{0}\rvert^{2}q_{0}\right)=0,$
@@
$\omega_{0}^{2}=\frac{\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})}{1+i\beta{c_{r}}^{2}\frac{\lvert q_{0}\rvert^{2}}{\lvert\operatorname{{\mathcal{D}}}\rvert^{2}}}.$ (19)
@@ 5 Numerical experiments: A nonlinear system of two resonators
To illustrate the results from Section 4, some numerical results will be presented in this section. Concretely, we will study the nonlinearity $i\beta\omega^{2}\lvert u\rvert^{2}u$ for the Helmholtz system (6), as discussed in Remark 4.8. The describing equation for the leading-order behavior of the subwavelength modes, is in that case given by
$\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})q_{0}-\omega_{0}^{2}\left(q_{0}+i\beta{c_{r}}^{2}(V^{vol})^{2}\lvert q_{0}\rvert^{2}q_{0}\right)=0,$ (20)
where the entries of the vector $q_{0}$ correspond to the average of the subwavelength resonant solutions and $\lvert q_{0}\rvert^{2}q_{0}$ is to be understood to act component-wise on the entries of $q_{0}$.
In the case where $\operatorname{{\mathcal{D}}}=B_{1}\cup B_{2}$ and $B_{1}$ and $B_{2}$ satisfy certain symmetry conditions, the resonant solutions to the linear system (2) remain resonant solutions of the nonlinear system.
@@
$v_{0}=\begin{pmatrix}a\\ a\end{pmatrix}\text{ and }v_{1}=\begin{pmatrix}-a\\ a\end{pmatrix}$
are solutions to equation (20) with
$\omega_{0}=\frac{\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{1,1}+\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{2,2}}{\left(1+i\beta\frac{{c_{r}}^{2}}{\lvert B_{1}\rvert^{2}}\lvert a\rvert^{2}\right)}\quad\text{ and }\quad\omega_{1}=\frac{\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{1,1}-\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{2,2}}{\left(1+i\beta\frac{{c_{r}}^{2}}{\lvert B_{1}\rvert^{2}}\lvert a\rvert^{2}\right)},$ (21)
respectively.
@@
with $C_{1},C_{2}\in\mathbb{R}$. Furthermore, $V(1,1)=V(2,2)=\lvert B_{1}\rvert^{-1}$. Thus, a basis of eigenvectors of $V\operatorname{cap}(\operatorname{{\mathcal{D}}})$ is given by
@@
$\begin{split}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})v_{j}-\omega_{0}^{2}\left(v_{j}+i\beta{c_{r}}^{2}(V^{vol})^{2}\lvert a\rvert^{2}v_{j}\right)\\ =\left(\lambda_{j}-\omega_{0}^{2}\left(1+i\beta\frac{{c_{r}}^{2}}{\lvert B_{1}\rvert^{2}}\lvert a\rvert^{2}\right)\right)v_{j}.\end{split}$
@@
$(\omega_{j}(a))^{2}=\frac{\lambda_{j}}{\left(1+i\beta\frac{{c_{r}}^{2}}{\lvert B_{1}\rvert^{2}}\lvert a\rvert^{2}\right)}$
@@
In the symmetric dimer case, the resonant solutions to the linear system are preserved for all amplitudes of the nonlinearity given in Remark 4.8. Comparing equations (19) and (21), one can observe that the dependence of the subwavelength resonances $\omega(a)$ on the amplitude $a$ is reminiscent of the monomer case discussed in Remark 4.8.
Of course, in the nonlinear setting with a cubic nonlinearity, one expects a third solution. For this solution to occur, the magnitude of the solution modes needs to be sufficiently large. In Figure 2(a), the solutions $q_{0}$ are displayed for a domain $\operatorname{{\mathcal{D}}}$ in the setting of Lemma 5.1. Concretely, the solutions of equation (20) are computed for $B_{1}=B_{0.2}(-\frac{1}{2}e_{3})$ and $B_{2}=B_{0.2}(\frac{1}{2}e_{3})$. The material parameters are chosen as follows ${c_{r}}=1$ and $\beta=0.1$.
Due to the complex nature of the solutions we decided to plot the magnitude of each entry $\lvert(q_{0})_{1}\rvert$, $\lvert(q_{0})_{2}\rvert$ and display the phase of $(q_{0})_{1}/(q_{0})_{2}$ as the respective color of the curve.
One can see the solutions predicted from Lemma 5.1 as the diagonal line in Figure 1(a) (the predicted solutions lie on top of each other). For a sufficiently big magnitude, one sees another solution curve occurring. The amplitudes of the entries $\lvert(q_{0})_{1}\rvert$ and $\lvert(q_{0})_{2}\rvert$ are symmetric and the phase difference is conjugate, as one would expect from a symmetric system. This is made precise in the following lemma.
@@
Let $\operatorname{{\mathcal{D}}}$ be a smooth bounded domain with well-separated connected components $B_{1},~{}B_{2}$ with connected boundaries. Suppose that $\operatorname{{\mathcal{D}}}=B_{1}\cup B_{2}$ is preserved under the mirror symmetry $P(x_{1},x_{2},x_{3})=(x_{1},x_{2},-x_{3})$ and $P(B_{1})=P(B_{2})$. Then, if $(\omega,q_{0})$ is a solution of (20), then
$\left(\omega,p_{0}\right)\quad\text{ with }\quad p_{0}=\begin{pmatrix}(q_{0})_{2}\\ (q_{0})_{1}\end{pmatrix}$
@@
When $\operatorname{{\mathcal{D}}}$ satisfies the above symmetry, it follows that $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ is symmetric. Plugging $(\omega,p_{0})$ into equation (20), one observes that the first component of (20) is given by the second component when plugging in $(\omega,q_{0})$ and the second is equal to the first. Thus, both are equal to zero and $(\omega,p_{0})$ is a solution to (20). ∎

### v3

@@ 4 Nonlinear subwavelength resonances
It is important to note that the techniques in this section are not limited to $g(u)=\lvert\frac{\partial u}{\partial t}\rvert^{2}\frac{\partial u}{\partial t}$, but can be used for any $g(u)$ that satisfies that $g(\exp(i\omega t)u(x))$ for $u\in H^{1}(\operatorname{{\mathcal{D}}})$ is of order $O(\omega^{2})$ as $\omega\rightarrow 0$ and that satisfies a gauge invariance in the sense that
$g(\exp(i\omega t)u(x))=\exp(i\omega t)v(x),$
for some $v\in H^{1}(\operatorname{{\mathcal{D}}})$ that depends on $u$ and $\omega$ but is constant with respect to $t$.
@@
$\left\{\begin{array}[]{rll}\displaystyle\Delta u+\frac{\omega^{2}}{{c_{r}}^{2}}u&={-}i\beta\lvert\omega\rvert^{2}\omega\lvert u\rvert^{2}u&\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial u}{\partial\nu}&=\delta\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u]&\text{on }\partial\operatorname{{\mathcal{D}}},\end{array}\right.$
@@
4.1 Variational characterization of nonlinear resonances
The following semilinear form will play the role of $a_{\omega,\delta}$ in the nonlinear setting.
@@ 4.1 Variational characterization of nonlinear resonances
Suppose that the smooth, bounded domain $\operatorname{{\mathcal{D}}}\in\operatorname{{\mathbb{R}^{3}}}3$ has well-separated connected components $B_{1},\ldots,B_{N}$ with connected boundaries. Define the semilinear form $\tilde{a}_{\omega,\delta}:H^{1}(\operatorname{{\mathcal{D}}})\times H^{1}(\operatorname{{\mathcal{D}}})\rightarrow\mathbb{C}$ associated to $\operatorname{{\mathcal{D}}}$ by
@@
Recall here from the Gagliardo-Nirenberg-Sobolev inequalities that $H^{1}(\operatorname{{\mathcal{D}}})\subset L^{4}(\operatorname{{\mathcal{D}}})$. With the form $\tilde{a}_{\omega,\delta}$ it is possible to characterize the resonances of the nonlinear system (6) in a similar fashion as in the linear case in Theorem 3.5 and Corollary 3.6. To this end, let $\beta$ be fixed. For $f\in L^{\infty}(\operatorname{{\mathcal{D}}})$, we can obtain from [27] (see also [9, 14]) that there exists a neighborhood $U$ of $0\in\mathbb{R}$ and a neighborhood $V$ of $0\in\mathbb{C}$ such that for all $\delta\in U$ and for all $\omega\in V$, the following variational problem:
@@
has a unique solution $u_{f}(\omega,\delta)$.
With this property at hand, the following variational characterization of small resonances of the nonlinear system (6) is possible.
@@
i.e., with $f=\sum_{i=1}^{N}a_{i}\int_{B_{i}}1_{B_{j}}$, satisfies
$\int_{B_{i}}u_{f}\operatorname{dvol}=a_{i},\quad\forall i\in\{1,\ldots,N\}.$ (16)
@@
4.2 Asymptotics of nonlinear subwavelength resonances
In this subsection, we will determine the solution of the variational problem (15) asymptotically and then derive asymptotics for subwavelength resonances of the nonlinear system (6). To this end, we consider the PDE formulation of the variational problem (15)
@@ 4.2 Asymptotics of nonlinear subwavelength resonances
Inserting this into (17) and identifying powers of $\delta$ and $\omega$, we obtain the following cascade of equations characterizing the functions $(u_{p,k,l})_{p,k\geq 0}$:
@@
This proposition allows us to derive the following fundamental theorem of nonlinear subwavelength physics by deriving the asymptotics of the nonlinear subwavelength resonances in terms of $\delta$.
@@
Let $\{q_{0},v_{2},\ldots,v_{N}\}$ be a basis of eigenvectors of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$, let $V^{vol}$ be as in Definition 2.3 and denote by $\Pi:\mathbb{C}^{N}\rightarrow\mathbb{C}q_{0}$ the projection along $\operatorname{span}(v_{2},\ldots,v_{N})$ onto $\mathbb{C}q_{0}$. Then the subwavelength resonance $\omega(\delta)=\omega_{0}\sqrt{\delta}+\omega_{1}\delta+O(\delta^{\frac{3}{2}})$ of (6) associated to $q_{0}$ satisfies the following asymptotics
$\omega_{0}\text{ is an eigenvalue of }\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\text{ with eigenvector }q_{0},$
$\omega_{1}=\frac{q_{0}\cdot\Pi\left[\frac{-i}{4\pi{c_{0}}}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}})q_{0}-\lvert\omega_{0}\rvert^{2}i\beta{c_{r}}^{2}(V^{vol})^{2}\lvert q_{0}\rvert^{2}q_{0}\right]}{2\lVert q_{0}\rVert^{2}},$ (18)
where $\lvert q_{0}\rvert^{2}q_{0}$ is to be understood component-wise.
@@
where $\lvert q(\delta)\rvert^{2}q(\delta)$ is to be understood component-wise. Making the Ansatz $\omega(\delta)=\omega_{0}\sqrt{\delta}+\omega_{1}\delta+O(\delta^{\frac{3}{2}})$ and $q(\delta)=q_{0}+q_{1}\sqrt{\delta}+O(\delta)$ and inserting it into the above equation leads to the following cascade of equations:
@@
It then follows that $\omega_{0}$ is an eigenvalue of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ with eigenvector $q_{0}$ and $\omega_{1}$ is given by
@@
Note that (19) has some similarities with the discrete Gross-Pitaevskii and Aubry-André-Harper models, which are used as discrete approximations of nonlinear Schrödinger equations (with cubic nonlinearities) in quantum mechanics; see, e.g., [10, 23, 18]. Such quantum mechanical models are typically based on tight-binding approximations and include only nearest-neighbour interaction terms (long-range interactions are neglected). They correspond to (19) with $\operatorname{cap}^{gen}$ being tridiagonal [13]. We refer to [26, 17, 15, 21, 20] for their analysis.
@@
As in [1], if we consider the scattering of an incident wave $u_{in}$ by a system of nonlinear subwavelength resonators, then based on Theorem 4.5, we can approximate the scattered field $u-u_{in}$ by a few subwavelength modes involving only the nonlinear subwavelength resonances and their associated eigenmodes, so as to retain the physically derived coupling between resonators.
As already mentioned at the beginning of this section, the above procedure is not limited to $g(u)=\lvert\frac{\partial u}{\partial t}\rvert^{2}\frac{\partial u}{\partial t}$ but solely relies on the property that $g(u)$ is gauge invariant and $g(u(x)\exp(i\omega t))$ is of order $O(\omega^{2})$ as $\omega\rightarrow 0$. The following remark is an illustration of this fact if the order is $O(\omega^{2})$ as $\omega\rightarrow 0$.
@@
If one changes the nonlinearity $i\beta\lvert\omega\rvert^{2}\omega\lvert u\rvert^{2}u$ to $-\beta\omega^{2}\lvert u\rvert^{2}u$, the nonlinearity is of order $\omega^{2}$ thus influencing the leading-order behavior of the subwavelength resonances. To be precise, the cascading system of equations from the proof of Proposition 4.4 is then given by
$-\Delta u_{p,k,l}+\sum_{i=1}^{N}\left(\int_{B_{i}}u_{p,k,l}\operatorname{dvol}\right)1_{B_{i}}$ $=\frac{1}{{c_{r}}^{2}}u_{p-2,k,l}-\beta\left(\lvert u_{f}\rvert^{2}u_{f}\right)_{p-2,k,l}+\sum_{j=1}^{N}a_{j}1_{B_{j}}\delta_{p=0}\delta_{k=0}\delta_{l=0},$
@@
$\begin{split}u_{f}(\omega,\delta)=\sum_{j=1}^{N}\frac{a_{j}}{\lvert B_{j}\rvert}1_{B_{j}}+\omega^{2}\left(\sum_{j=1}^{N}\frac{a_{j}}{{c_{r}}^{2}\lvert B_{j}\rvert^{2}}1_{B_{j}}-\beta\sum_{j=1}^{N}\frac{\lvert a_{j}\rvert^{2}a_{j}}{\lvert B_{j}\rvert^{4}}1_{B_{j}}\right)\quad\quad\quad\quad\quad\quad\\ +\delta\sum_{j=1}^{N}a_{j}\left(-\sum_{l=1}^{N}\frac{\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}}{\lvert B_{j}\rvert\lvert B_{l}\rvert^{2}}1_{B_{l}}+\tilde{u}^{j}_{0,1}\right)\quad\quad\quad\quad\\ +\omega\delta\sum_{j=1}^{N}a_{j}\left(\sum_{l=1}^{N}\left(\frac{i}{4\pi\lvert B_{j}\rvert\lvert B_{l}\rvert^{2}{c_{0}}}(\operatorname{cap}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}}))_{l,j}\right)1_{B_{l}}+\tilde{u}^{j}_{1,1}\right)+O((\omega^{2}+\delta)^{2}).\end{split}$
In that case, the leading order asymptotics of the subwavelength resonances $\omega(\delta)=\omega_{0}\sqrt{\delta}+O(\delta)$ are characterized by the nonlinear eigenvalue problem
$\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})q_{0}-\omega_{0}^{2}\left(q_{0}-\beta{c_{r}}^{2}(V^{vol})^{2}\lvert q_{0}\rvert^{2}q_{0}\right)=0,$ (20)
@@
$\omega_{0}^{2}=\frac{\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})}{1-\beta{c_{r}}^{2}\frac{\lvert q_{0}\rvert^{2}}{\lvert\operatorname{{\mathcal{D}}}\rvert^{2}}}.$ (21)
@@ 5 Numerical experiments: A nonlinear system of two resonators
To illustrate the results from Section 4, some numerical results will be presented in this section. Concretely, we will study the nonlinearity $-\beta\omega^{2}\lvert u\rvert^{2}u$ for the Helmholtz system (6), as discussed in Remark 4.8. The describing equation for the leading-order behavior of the subwavelength modes, is in that case given by
$\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})q_{0}-\omega_{0}^{2}\left(q_{0}-\beta{c_{r}}^{2}(V^{vol})^{2}\lvert q_{0}\rvert^{2}q_{0}\right)=0,$ (22)
where the entries of the vector $q_{0}$ correspond to the average of the subwavelength resonant solutions inside the connected components $B_{1},\ldots,B_{N}$ and $\lvert q_{0}\rvert^{2}q_{0}$ is to be understood to act component-wise on the entries of $q_{0}$.
In the case where $\operatorname{{\mathcal{D}}}=B_{1}\cup B_{2}$ and $B_{1}$ and $B_{2}$ satisfy certain symmetry conditions, the leading order behavior of the resonant solutions to the linear system (2) is not influenced by the nonlinearity, however the leading order behavior of the resonances changes.
@@
$v_{0}=\begin{pmatrix}a\\ a\end{pmatrix}\text{ and }~{}v_{1}=\begin{pmatrix}-a\\ a\end{pmatrix}$
with
$\omega_{0}=\frac{\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{1,1}+\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{2,2}}{\left(1-\beta\frac{{c_{r}}^{2}}{\lvert B_{1}\rvert^{2}}\lvert a\rvert^{2}\right)}\quad\text{ and }\quad\omega_{1}=\frac{\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{1,1}-\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{2,2}}{\left(1-\beta\frac{{c_{r}}^{2}}{\lvert B_{1}\rvert^{2}}\lvert a\rvert^{2}\right)},$ (23)
respectively, are solutions to equation (22).
@@
with $C_{1},C_{2}\in\mathbb{R}$. Furthermore, $(V^{vol})_{1,1}=(V^{vol})_{2,2}=\lvert B_{1}\rvert^{-1}=\lvert B_{2}\rvert^{-1}$. Thus, a basis of eigenvectors of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})={c_{r}}^{2}V^{vol}\operatorname{cap}(\operatorname{{\mathcal{D}}})$ is given by
@@
$\begin{split}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})v_{j}-\omega_{0}^{2}\left(v_{j}-\beta{c_{r}}^{2}(V^{vol})^{2}\lvert a\rvert^{2}v_{j}\right)\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\\ =\left(\lambda_{j}-\omega_{0}^{2}\left(1-\beta\frac{{c_{r}}^{2}}{\lvert B_{1}\rvert^{2}}\lvert a\rvert^{2}\right)\right)v_{j},\quad\text{ for }j\in\{0,1\}.\end{split}$
@@
$(\omega_{j}(a))^{2}=\frac{\lambda_{j}}{\left(1-\beta\frac{{c_{r}}^{2}}{\lvert B_{1}\rvert^{2}}\lvert a\rvert^{2}\right)}$
@@
In the symmetric dimer case, the leading order asymptotics of the resonant solutions to the linear system are preserved for all amplitudes of the nonlinearity given in Remark 4.8. Comparing equations (21) and (23), one can observe that the dependence of the leading order of the subwavelength resonances $\omega(a)$ on the amplitude $a$ is reminiscent of the monomer case discussed in Remark 4.8.
Of course, in the nonlinear setting with a cubic nonlinearity, one expects a third solution. For this solution to occur, the magnitude of the solution modes needs to be sufficiently large. In Figure 2(a), the solutions $q_{0}$ to equation (16) are displayed for a domain $\operatorname{{\mathcal{D}}}$ in the context of Lemma 5.1. Concretely, the solutions $(q_{0},\omega_{0})$ of equation (22) are computed for $B_{r_{1}}=B_{0.2}(-\frac{1}{2}e_{3})$ and $B_{r_{2}}=B_{0.2}(\frac{1}{2}e_{3})$, where $B_{r}(x)$ for some $r>0$ and $x\in\mathbb{R}^{3}$ is defined as $B_{r}(x)=\{y\in\mathbb{R}^{3}:\lvert x-y\rvert<r\}$. The material parameters are chosen as follows ${c_{r}}=1$ and $\beta=-0.1i$.
Due to the complex nature of the solutions, we decided to plot the magnitude of each entry $\lvert(q_{0})_{1}\rvert$, $\lvert(q_{0})_{2}\rvert$ and display the phase of $(q_{0})_{1}/(q_{0})_{2}$ as the respective color of the curve.
The solutions predicted from Lemma 5.1 can be seen as the diagonal line in Figure 1(a) (the predicted solutions lie on top of each other). For a sufficiently big magnitude, one sees another solution curve occurring. The amplitudes of the entries $\lvert(q_{0})_{1}\rvert$ and $\lvert(q_{0})_{2}\rvert$ are symmetric and the phase difference is conjugate, as expected from a symmetric system. This is made more precise in the following lemma.
Figure 5.1: In these plots the solution curves of equation (22) for different choices of $\operatorname{{\mathcal{D}}}$ are plotted. From left to right, $\operatorname{{\mathcal{D}}}=B_{0.2}\left(-\frac{1}{2}e_{3}\right)\cup B_{0.2}\left(\frac{1}{2}e_{3}\right)$, $\operatorname{{\mathcal{D}}}=B_{0.2}\left(-\frac{1}{2}e_{3}\right)\cup B_{0.21}\left(\frac{1}{2}e_{3}\right)$ and $\operatorname{{\mathcal{D}}}=B_{0.2}\left(-\frac{1}{2}e_{3}\right)\cup B_{0.22}\left(\frac{1}{2}e_{3}\right)$, respectively. The remaining parameters are given by $\beta=-0.1i$ and ${c_{r}}=1$. The colorful curves trace the magnitude of the entries of different solutions $q_{0}$ to equation (22) (the respective $\omega_{0}$ can be found in Figure 5.2). The colors account for the complex valued solutions $q_{0}\in\mathbb{C}^{2}$ and depict the phase of $(q_{0})_{1}/(q_{0})_{2}$. The black dashed curves are the solutions to the respective linear problem when $\beta=0$.
@@
Let $\operatorname{{\mathcal{D}}}$ be a smooth bounded domain with well-separated connected components $B_{1},~{}B_{2}$ with connected boundaries. Suppose that $\operatorname{{\mathcal{D}}}=B_{1}\cup B_{2}$ is preserved under the mirror symmetry $P(x_{1},x_{2},x_{3})=(x_{1},x_{2},-x_{3})$ and $P(B_{1})=P(B_{2})$. Then, if $(\omega_{0},q_{0})$ is a solution of (22), then
$\left(\omega_{0},p_{0}\right)\quad\text{ with }\quad p_{0}=\begin{pmatrix}(q_{0})_{2}\\ (q_{0})_{1}\end{pmatrix}$
@@
When $\operatorname{{\mathcal{D}}}$ satisfies the above symmetry, it follows that $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ is symmetric. Plugging $(\omega_{0},p_{0})$ into equation (22), one observes that the first component of (22) is given by the second component when plugging in $(\omega_{0},q_{0})$ and the second is equal to the first. Thus, both are equal to zero and $(\omega_{0},p_{0})$ is a solution to (22). ∎